# Standard Lyndon loop words: weighted orders

### v1

@@ A.1. Examples
The first version of our code did not use the key results (Theorem 3.6 and Proposition 3.8), but was rather based on Remark 3.10, which is a simple generalization of [16, Proposition 2.26]. Thus, when evaluating $\ell(\alpha,d)$, the code simply goes through all the ways to split $\alpha$ into an ordered sum of simple roots, and distribute $d$ between the exponents of these simple roots while satisfying (3.12). In the tables below, we present examples of standard Lyndon loop words computed through this code (which also nicely illustrate the results of Theorem 3.6 and Proposition 3.8).
@@
Let us clarify the conventions in these tables:
  • $\bullet$
    In the column “Order”, the elements $i\in I$ are listed in the increasing order,
  • $\bullet$
    In the column “Weights”, the weights $c_{i}$ are listed with $i$ ordered as in [23].
In all the examples above, we choose to consider only the highest root $\theta$.
@@ A.2. The code
The second version of our code was written using Proposition 3.8, which provides an inductive way to compute exponents of $\ell(\alpha,d)$, thus drastically improving the code performance and allowing us to compute words for much larger values of the degree $d$ and the weights $c_{i}$. This code can be used at the following clickable link: [^2] The user should press the “Run” button and they will see the instructions and a small example afterwards. Type in the input in the console afterwards, following the instructions. Names of Lie algebra types for input are: A, B, C, D, G2, F4, E6, E7, E8. This code was written using C++23.
We have checked on many examples that both codes produce the same answers (while the second is much faster). The interested reader can use this code to check the results of this note as well as to compute standard Lyndon loop words.

### v2

@@ A.1. Examples
The first version of our code did not use the key results (Theorem 3.6 and Proposition 3.8), but was rather based on Remark 3.10, which is a simple generalization of [16, Proposition 2.26]. Thus, when evaluating $\ell(\alpha,d)$, the code simply goes through all the ways to split $\alpha$ into an ordered sum of simple roots, and distribute $d$ between the exponents of these simple roots while satisfying (3.12). In the table below, we present examples of standard Lyndon loop words computed through this code (which also nicely illustrate the results of Theorem 3.6 and Proposition 3.8).
@@
Let us clarify the conventions in this table:
  • $\bullet$
    In the column “Order”, the elements $i\in I$ are listed in the increasing order;
  • $\bullet$
    In the column “Weights”, the weights $c_{i}$ are listed with $i$ ordered as in [23];
  • $\bullet$
    In all these examples, we choose to consider only the highest root $\alpha=\theta$.
Let us also provide examples of standard Lyndon loop words for the remaining exceptional types (these were evaluated using our second code presented below):
@@
$E_{7}$ 1234567 4 5 3 7 3 2 5 25 $[4^{(3)}5^{(1)}6^{(0)}3^{(1)}7^{(2)}4^{(2)}2^{(2)}1^{(1)}3^{(1)}2^{(2)}4^{(3)}5^{(1)}6^{(0)}3^{(1)}7^{(2)}4^{(2)}5^{(1)}]$
$E_{7}$ 1234567 4 5 3 7 3 2 5 26 $[4^{(3)}5^{(1)}3^{(1)}7^{(2)}4^{(2)}2^{(2)}1^{(1)}3^{(1)}2^{(2)}4^{(3)}5^{(1)}6^{(0)}3^{(1)}7^{(2)}4^{(3)}5^{(1)}6^{(0)}]$
$E_{8}$ 14572386 1 32 13 3 10 9 6 15 46 $[8^{(3)}5^{(1)}4^{(0)}6^{(1)}5^{(1)}3^{(2)}4^{(0)}7^{(1)}6^{(1)}5^{(1)}2^{(6)}1^{(0)}3^{(2)}4^{(0)}2^{(6)}3^{(2)}$
$8^{(3)}5^{(1)}4^{(0)}6^{(1)}5^{(1)}3^{(2)}4^{(0)}7^{(1)}6^{(1)}5^{(1)}8^{(2)}2^{(6)}1^{(0)}]$
$E_{8}$ 14572386 1 32 13 3 10 9 6 15 47 $[8^{(3)}5^{(1)}4^{(0)}6^{(1)}5^{(1)}3^{(2)}7^{(1)}6^{(1)}2^{(6)}1^{(0)}8^{(3)}5^{(1)}4^{(0)}6^{(1)}5^{(1)}3^{(2)}$
$4^{(0)}7^{(1)}2^{(6)}3^{(2)}8^{(3)}5^{(1)}4^{(0)}6^{(1)}5^{(1)}3^{(2)}4^{(0)}2^{(6)}1^{(0)}]$
@@ A.2. The code
The second version of our code was written using Proposition 3.2 as well as Proposition 3.8 which provides an inductive way to compute exponents of $\ell(\alpha,d)$. This drastically improves the code performance, allowing us to compute words for much larger values of the degree $d$ and the weights $c_{i}$. This code can be used at the following clickable link (the interested reader can use this code to check the results of this note as well as to compute standard Lyndon loop words): [^2] The user should press the “Run” button and they will see the instructions and a small example afterwards. Type in the input in the console afterwards, following the instructions. Names of Lie algebra types for input are: A, B, C, D, G2, F4, E6, E7, E8. This code was written using C++23.